# Conghui He4, Xiaoyang Wang6, Fan Wu1

### v1

@@ 4.2 Video
Recent methods introduce architectures capable of isolating region-level tokens and maintaining object-specific trajectories over time. To capture fine-grained spatial and temporal details that are often ignored by former models , VideoRefer yuan2025videorefer employs a spatial–temporal object encoder that couples a spatial token extractor with adaptive temporal token merging, producing precise region representations for single-frame and multi-frame reasoning. To enhance video analysis precision and efficiency de2024video , which employs YOLO-V8 to detect objects and segment event-relevant regions, while Gemini-2.5-flash serves as a multimodal language backbone to transform these visual detections into structured textual reports. The framework introduces an incident–context fusion pipeline that aggregates temporal object trajectories and scene cues before passing them into Gemini, enabling the model to generate summaries that retain both fine-grained entity information and broader situational context. By coupling high-precision, anchor-free detection with multimodal reasoning, the system reduces dependence on exhaustive manual review and ensures that critical events are captured without sacrificing efficiency. sengonul2025abnormal offers a pipeline for abnormal event detection in surveillance that spans three stages: (1) preprocessing through background subtraction and noise reduction, (2) feature learning via an LSTM autoencoder modeling normal behavior trajectories, and (3) anomaly detection using distribution-based regularity scoring enhanced by local minima analysis for confinement of anomalous instances while achieving high AUCs on CUHK Avenue and UMN datasets.
@@ 4.3 Documents
Takeaway: Effective document understanding has shifted from unimodal text processing to sophisticated multimodal architectures that integrate textual, spatial, and visual information. Retrieval-Augmented Generation (RAG) and synthetic data are becoming crucial for enhancing factual accuracy and model robustness.
@@
TABLE III: Challenges of Unstructured Document Analysis.
Document Type Key Challenges / Required Capabilities
Short-form Documents
(e.g., Invoices, web pages)
Factual QA, concise summarization, key-value extraction.
Long-form Documents
(e.g., Scientific papers, reports)
Integrative comprehension, cross-document synthesis, handling distractors.
Multilingual Corpora
(e.g., Parallel/comparable documents)
High-fidelity translation, cross-lingual QA and summarization.
Core Challenges in Document Understanding. Document understanding is complicated. As shown in Table III, for short-form documents (e.g., invoices, web pages), a key bottleneck is form understanding jaume2019funsd , which involves tasks like accurate factual question answering by locating specific information amidst layout noise, generating concise summaries, and performing key-value extraction from structured content. For long-form documents (e.g., scientific papers, technical reports), the challenges escalate to integrative comprehension and cross-document synthesis, along with a robust capability for handling distractors. Finally, handling multilingual corpora requires not only high-fidelity translation but also complex downstream tasks like cross-lingual QA or summarization.
To tackle these challenges, the research community has developed several distinct technical paradigms. We categorize them into three primary approaches: (1) architecture designs ranging from unimodal to multimodal designs, with particular emphasis on fusion strategies and cross-modal interactions, (2) retrieval-augmented methods for enhancing factual accuracy and context, and (3) synthetic data generation techniques to address data scarcity and improve model training.
TABLE IV: Taxonomy of Multimodal Document Understanding Architectures.
Fusion Stage Architecture Type Representative Models Technical Characteristics
@@
CNN + Transformer Traditional pipelines Visual CNN, text Transformer, early concat
Intermediate Fusion Dual-stream Transformer LayoutLMv2, DocFormer Parallel text/image streams, cross-attention
@@
Encoder-Decoder OmniParser, CREPE Shared encoder, task-specific decoders
@@
Pipeline Integration OCR+Layout systems Independent modules, result combination
@@
Architecture Design for Document Understanding. As shown in Table IV, the evolution of document understanding architectures reflects a fundamental shift from processing individual modalities in isolation to sophisticated multimodal integration. Early approaches treated text, layout, and visual elements as separate channels, which often led to suboptimal performance due to the loss of critical cross-modal relationships inherent in document structure. Unlike plain text processing, document understanding inherently requires the simultaneous consideration of textual content, spatial layout, and visual appearance. This realization has driven the field toward multimodal architectures that can capture these complex interdependencies.
@@ 4.4 Program Language Analysis
Takeaway: The core of LLM-based program analysis is advanced data synthesis techniques. By creating high-quality “code-task pairs” through methods like autoformalization, iterative refinement, and RAG, models can tackle complex tasks from automated theorem proving to repository-level code completion.
In this section we introduce LLMs for program analysis, focusing on code-task pairs. Unlike raw source code, this data explicitly links program artifacts to their semantic properties or desired outputs, forming the basis for sophisticated analysis tasks. Examples include code paired with vulnerability labels, formal proofs, or natural language summaries. We first survey the downstream applications that this data enables and then describe the core methodologies for synthesizing this specialized data. Table V summarizes the core applications and the synthesis techniques that enable them.
TABLE V: Core Techniques in LLM-Based Program Analysis.
@@
ATP Task-Specific Generation
Iterative Refinement
Autoformalization
Expert Iteration
Vulnerability Detection Task-Specific Generation
Retrieval Prompting
Analysis Augmentation
In-Context Learning (ICL)
Code Summarization Few-Shot Learning
RAG
Synthesized Data Pairs
Example Retrieval
Repo-Level Completion RAG Adaptive/Iterative Retrieval
Context Fusion
@@
Program Vulnerability Detection. LLMs enhance software vulnerability detection by learning from rich program data and leveraging contextual reasoning. One direction focuses on program-analysis–based training, where models such as PDBERT (liu2024pdbert, ) are fine-tuned on code augmented with program analysis information—such as control and data dependencies—to enable fine-grained vulnerability classification. Related methods also incorporate decomposed control flow graphs (CFGs) to capture intra- and inter-path features for more precise vulnerability prediction zhang2023bulnerability . Complementarily, case-driven prompt engineering exploits the in-context learning capability of LLMs by retrieving similar code snippets along with their known vulnerabilities and feeding this context, together with a model-generated interpretation of the target code, back into the LLM to boost detection accuracy. Approaches like VUL-GPT liu2023gptvulnerability and other prompt-based systems zhou2024llmvulnerability exemplify this strategy, demonstrating notable improvements in identifying subtle security flaws.
@@
Takeaway: Reformulating 3D models using the 3D encoder and projector aligns the 3D space with the semantic space, enabling llms to understand spatial features. Using domain-specialized interpreters or cross-modal feature extractor can improve the efficiency of 3D-llms.
@@
Traditional 3D model analysis involves using 3D modeling software (e.g. Blender, 3ds Max, Maya) and need human input for manipulation, annotation, and interpretation. As LLMs evolve, their integration with 3D spatial data enables unprecedented capabilities in understanding and interacting with physical environments. This section provides a comprehensive overview of methodologies that allow LLMs to process, interpret, and generate 3D data from point clouds to Neural Radiance Fields (NeRFs), and examines their applications in tasks such as scene understanding, captioning, question-answering, and dialogue. We further discuss the role of LLM-powered agents in spatial reasoning, planning, and navigation, highlighting their potential to bridge natural language with the inherent complexities of 3D spaces. The fusion of LLMs with 3D data primarily manifests in three key areas: 3D Language Fusion (see Figure 9), which aligns 3D geometry with language descriptions; 3D-Derived Task Enhancement (see Figure 10), which leverages 3D information to improve performance on downstream tasks; and Cross-Modal Capability Refinement (see Figure 11), which uses existed experience on models (like 2-D image models) to refine 3D model’s understanding and generation abilities across different modalities.
TABLE VI: Overview of 3D-Language Fusion and Related Techniques.
@@
Figure 9: Overview of 3D Model Analysis with 3D Language Fusion.
@@
Figure 10: Overview of 3D Model Analysis with 3D-Derived Task Enhancement.
3D-Derived Task Enhancement Unlike 3D-Language Fusion which focuses on aligning spatial features with textual semantics, 3D-derived task enhancement emphasizes leveraging structural representations to improve downstream tasks that inherently require geometric reasoning (e.g., spatial question answering) as shown in Figure 10. The central challenge lies in (1) achieving genuine structural understanding rather than relying on superficial correlations, and (2) ensuring scalability, since incorporating high-dimensional 3D signals into large models can quickly lead to prohibitive memory and computation costs.
@@
Figure 11: Overview of 3D Model Analysis with Cross-Modal Capability Refinement.
@@
To handle the limitation of domain adaptation technique in learning discriminative features for target samples due to the lack of label information in target domain, Song et al. song2023self improve image-mesh retrieval by combining adversarial domain alignment with self-supervised contrastive learning on unlabeled 3D data. The objective integrates (1) supervised discriminative learning on labeled 2D images, (2) contrastive losses on multi-view 3D renderings that pull views of the same object together and push different objects apart, and (3) adversarial terms to minimize domain gaps between 2D and 3D embeddings. An entropy-aware memory bank maintains representative negative samples and reduces false negatives, producing highly discriminative 3D embeddings and boosting retrieval accuracy.
To equip multimodal LLMs with 3D perception from single images, LLMI3D yang2024llmi3d augments a pretrained MLLM using parameter-efficient fine-tuning (PEFT) and three spatial reasoning modules: (1) Spatial-Enhanced Local Feature Mining sharpens object boundary and micro-geometry features via multi-scale attention with positional cues, (2) 3D Query Token-Derived Info Decoding introduces explicit geometric-query tokens with regression heads for precise numeric output, and (3) Geometry Projection-based Reasoning corrects for focal-length and perspective distortion by projecting features through differentiable camera models. The result shows that carefully designed spatial modules with PEFT can make general MLLMs competitive for 3D geometric reasoning without full-model retraining.
TABLE VII: Comparison of 3D Generation Methods across different paradigms.
@@
3D Data for LLM. Large language models have shown remarkable capabilities in reasoning over text and 2D imagery, but extending them to 3D domains introduces unique challenges. Unlike text or images, 3D data inherently encodes geometry, spatial relations, and multi-view consistency, which makes it both richer in information and more difficult to model. Recent research has therefore explored different paradigms for leveraging or directly generating 3D representations conditioned on language, aiming to unlock applications such as text-to-3D asset creation, embodied AI, and interactive design. Table VII Existing methods can be broadly grouped into three categories: optimization-based approaches that adapt powerful 2D diffusion priors for 3D generation, feed-forward pipelines that prioritize speed and determinism, and emerging native 3D generative models that attempt to learn 3D distributions directly.
@@ 5 Heterogeneous Data Analysis
Takeaway: Heterogeneous data analysis requires capturing both structural and semantic information across diverse modalities, a task that traditional human-driven approaches handle inefficiently. LLM- and VLM-powered agents now enable cross-modal alignment, retrieval, and reasoning through natural language, significantly reducing manual effort and supporting more intelligent, context-aware analysis.
@@
XMODE nooralahzadeh2024explainable assumes that heterogeneous metadata is organized within a database and leverages LLMs to decompose natural language queries into subtasks such as text-to-SQL generation and image analysis. For instance, given the query “Plot the number of paintings that depict war in each century”, XMODE generates an operation sequence comprising: (1) retrieving painting metadata from the database via SQL, (2) applying VLMs to determine whether each image depicts war, and (3) producing visualization code. In practice, however, heterogeneous data is often disorganized. To address this challenge, Wang et al. wang2024interactive encode data objects using modality-specific models (e.g., BERT for text, CLIP for images) to construct an index. They further introduce a vector weight learning model that adaptively adjusts modality weights to better capture their relative importance in similarity measurement wang2024must . The user query is used to retrieve relevant data through semantic similarity for final answer generation.
@@ 6.1 Structured Data
Complex Analysis. Structured data encapsulates both semantic information (e.g., cell contents in tables) and structural information (e.g., key–value relationships in tables or edge connections in graphs). Effective analysis requires integrating these two dimensions. For simple tasks, LLMs can often derive answers directly from explicit semantic cues or straightforward structural patterns. However, complex tasks demand deeper comprehension and multi-hop reasoning across semantics and structure, which remain challenging for current LLMs.
When LLMs alone fall short in handling such tasks end-to-end, multi-agent collaboration emerges as a promising direction. By decomposing a complex task into simpler subtasks, each agent can specialize in a narrower scope. Through prompt engineering or fine-tuning, LLMs can be tailored for specific subtasks, while intermediate outputs enhance both performance and interpretability.
Open World Task Adaptation. Structured data spans diverse domains with varying specifications and task requirements. LLMs trained for general-purpose analysis may fail to meet domain-specific needs. For instance, analyzing daily forms differs substantially from rigorous analyses in business or climate applications. Moreover, heterogeneity in system specifications, such as SQL dialects, poses additional adaptation challenges.
To address these issues, domain adaptation through task-specific training or fine-tuning is essential. As manual annotation is often insufficient to support large-scale training, efficient automatic data generation methods warrant further exploration. In addition, techniques such as automatic query conversion and intermediate representations are crucial for bridging modality gaps, particularly between natural language and diverse SQL dialects.

### v2

@@ 4.2 Video
To enhance video analysis precision and efficiency, de2024video uses YOLO-V8 for object detection and event-region segmentation, while Gemini-2.5-flash converts these visual cues into structured textual reports. An incident–context fusion pipeline aggregates temporal object trajectories and scene information before input to Gemini, producing summaries that preserve fine-grained entity details and overall context. By combining high-precision, anchor-free detection with multimodal reasoning, the framework reduces manual review while reliably capturing critical events. sengonul2025abnormal proposes a three-stage pipeline for surveillance anomaly detection: $(i)$ preprocessing with background subtraction and noise reduction, $(ii)$ feature learning via an LSTM autoencoder of normal behavior trajectories, and $(iii)$ anomaly detection using distribution-based regularity scoring with local minima analysis.
@@
Takeaway: Through reformulating videos as structured token sequences containing temporal and visual evidence, LLMs can track and understand specific events in object level. Complex Data Understanding (O1): Processing spatial semantics and temporal dynamics, multi-modal cues for emotional analysis, and 3D human behavior wang2024grounded ; weng2025artificial . NL-Based Interface (O2): Driven through natural language using methods including adapters for instruction-conditioned reasoning chen2024timemarker ; li2024llms . Semantic Operators (O3): Perform instruction-aware relative temporal localization to recognize conceptual relationships and contextual details huang2024image ; de2024video . Autonomous Evolution (O4): Advance through adaptive methods like dynamic frame sampling and token merging chen2024timemarker ; yuan2025videorefer .
@@ 4.3 Documents
Core Challenges in Document Understanding. Document understanding is complicated. For short-form documents (e.g., invoices, web pages), a key bottleneck is form understanding jaume2019funsd , which involves tasks like accurate factual question answering by locating specific information amidst layout noise, generating concise summaries, and performing key-value extraction from structured content. For long-form documents (e.g., scientific papers, technical reports), the challenges escalate to integrative comprehension and cross-document synthesis, along with a robust capability for handling distractors. Finally, handling multilingual corpora requires not only high-fidelity translation but also complex downstream tasks like cross-lingual QA or summarization.
To tackle these challenges, the research community has developed several distinct technical paradigms. We categorize them into three primary approaches: $(i)$ architecture designs ranging from unimodal to multimodal designs, with particular emphasis on fusion strategies and cross-modal interactions, $(ii)$ retrieval-augmented methods for enhancing factual accuracy and context, and $(iii)$ synthetic data generation techniques to address data scarcity and improve model training.
TABLE III: Taxonomy of Multimodal Document Understanding Architectures.
Fusion Stage Architecture Example Characteristics
@@
CNN + Transformer Pipelines Visual CNN, text Transformer, early concat
Intermediate Fusion Dual-stream Transformer LayoutLMv2 DocFormer Parallel text/image streams, cross-attention
@@
Encoder-Decoder OmniParser CREPE Shared encoder, task-specific decoders
@@
Pipeline Integration OCR+Layout Independent modules, result combination
@@
Architecture Design for Document Understanding. As shown in Table III, the evolution of document understanding architectures reflects a fundamental shift from processing individual modalities in isolation to sophisticated multimodal integration. Early approaches treated text, layout, and visual elements as separate channels, which often led to suboptimal performance due to the loss of critical cross-modal relationships inherent in document structure. Unlike plain text processing, document understanding inherently requires the simultaneous consideration of textual content, spatial layout, and visual appearance. This realization has driven the field toward multimodal architectures that can capture these complex interdependencies.
@@
Takeaway: LLM-based agents enhance document analysis by: Complex Data Understanding (O1): Leveraging multimodal architectures like LayoutLM (xu2020layoutlm, ) and DocLLM wang2024docllm to holistically integrate textual, spatial, and visual information from complex layouts. NL-Based Interface (O2): Enabling prompt-guided interaction where natural language queries direct visual analysis, as seen in methods like VisFocus (abramovich2024visfocus, ). Semantic Operators (O3): Employing advanced RAG systems to retrieve and synthesize information from both text and images, providing context-aware answers and summaries (lewis2020retrieval, ; lei2024visdommultimodal, ). Autonomous Evolution (O4): Utilizing iterative refinement and self-correction loops in synthetic data generation to progressively improve layout and content quality (shi2025layoutcot, ; zhang2024vascar, ).
@@ 4.4 Program Language Analysis
In this section we introduce LLMs for program analysis, focusing on code-task pairs. Unlike raw source code, this data explicitly links program artifacts to their semantic properties or desired outputs, forming the basis for sophisticated analysis tasks. Examples include code paired with vulnerability labels, formal proofs, or natural language summaries. We first survey the downstream applications that this data enables and then describe the core methodologies for synthesizing this specialized data. Table IV summarizes the core applications and the synthesis techniques that enable them.
TABLE IV: Core Techniques in LLM-Based Program Analysis.
@@
ATP Task-Specific Generation Iterative Refinement Autoformalization Expert Iteration
Code Summarization Few-Shot Learning RAG Synthesized Data Pairs Example Retrieval
Repo-Level Completion RAG Adaptive/Iterative Retrieval Context Fusion
@@
Takeaway: In program analysis, LLM-based agents demonstrate significant capabilities by: Complex Data Understanding (O1): Parsing entire code repositories with RAG techniques or translating natural language problems into formal code via autoformalization (wu2024repoformer, ; xin2024deepseek-prover, ). NL-Based Interface (O2): Automating the translation of requirements expressed in natural language into formal, machine-verifiable proofs or concise code summaries (xin2024deepseek-prover, ; geng2023llmfewshot, ). Semantic Operators (O3): Implementing adaptive retrieval mechanisms in repository-level code completion to find semantically relevant code snippets, going beyond simple keyword matching (zhang2023repocoder, ; liang2024repofuse, ). Autonomous Evolution (O4): Employing techniques like expert iteration, reinforcement learning with verifier feedback, and self-debugging to continuously refine their problem-solving and code generation abilities (xin2024deepseek-prover, ; chen2023teaching, ).
@@
Traditional 3D model analysis involves using 3D modeling software (e.g. Blender, 3ds Max, Maya) and need human input for manipulation, annotation, and interpretation. As LLMs evolve, their integration with 3D spatial data enables unprecedented capabilities in understanding and interacting with physical environments. This section provides a comprehensive overview of methodologies that allow LLMs to process, interpret, and generate 3D data from point clouds to Neural Radiance Fields (NeRFs), and examines their applications in tasks such as scene understanding, captioning, question-answering, and dialogue. We further discuss the role of LLM-powered agents in spatial reasoning, planning, and navigation, highlighting their potential to bridge natural language with the inherent complexities of 3D spaces. As shown in Figure 9, the fusion of LLMs with 3D data primarily manifests in three key areas: 3D Language Fusion, which aligns 3D geometry with language descriptions; 3D-Derived Task Enhancement, which leverages 3D information to improve performance on downstream tasks; and Cross-Modal Capability Refinement, which uses existed experience on models (like 2-D image models) to refine 3D model’s understanding and generation abilities across different modalities.
TABLE V: Overview of 3D-Language Fusion and Related Techniques.
@@
Figure 9: Overview of 3D Model Analysis.
@@
3D-Derived Task Enhancement Unlike 3D-Language Fusion which focuses on aligning spatial features with textual semantics, 3D-derived task enhancement emphasizes leveraging structural representations to improve downstream tasks that inherently require geometric reasoning (e.g., spatial question answering) as shown in Figure 9. The central challenge lies in $(i)$ achieving genuine structural understanding rather than relying on superficial correlations, and $(ii)$ ensuring scalability, since incorporating high-dimensional 3D signals into large models can quickly lead to prohibitive memory and computation costs.
@@
To handle the limitation of domain adaptation technique in learning discriminative features for target samples due to the lack of label information in target domain, Song et al. song2023self improve image-mesh retrieval by combining adversarial domain alignment with self-supervised contrastive learning on unlabeled 3D data. The objective integrates $(i)$ supervised discriminative learning on labeled 2D images, $(ii)$ contrastive losses on multi-view 3D renderings that pull views of the same object together and push different objects apart, and $(iii)$ adversarial terms to minimize domain gaps between 2D and 3D embeddings. An entropy-aware memory bank maintains representative negative samples and reduces false negatives, producing highly discriminative 3D embeddings and boosting retrieval accuracy.
To equip multimodal LLMs with 3D perception from single images, LLMI3D yang2024llmi3d augments a pretrained MLLM using parameter-efficient fine-tuning (PEFT) and three spatial reasoning modules including $(i)$ Spatial-Enhanced Local Feature Mining sharpens object boundary and micro-geometry features via multi-scale attention with positional cues, $(ii)$ 3D Query Token-Derived Info Decoding introduces explicit geometric-query tokens with regression heads for precise numeric output, and $(iii)$ Geometry Projection-based Reasoning corrects for focal-length and perspective distortion by projecting features through differentiable camera models. The result shows that carefully designed spatial modules with PEFT can make general MLLMs competitive for 3D geometric reasoning without full-model retraining.
TABLE VI: Comparison of 3D Generation Methods across different paradigms.
@@
3D Data for LLM. Large language models have shown remarkable capabilities in reasoning over text and 2D imagery, but extending them to 3D domains introduces unique challenges. Unlike text or images, 3D data inherently encodes geometry, spatial relations, and multi-view consistency, which makes it both richer in information and more difficult to model. Recent research has therefore explored different paradigms for leveraging or directly generating 3D representations conditioned on language, aiming to unlock applications such as text-to-3D asset creation, embodied AI, and interactive design. Table VI Existing methods can be broadly grouped into three categories: optimization-based approaches that adapt powerful 2D diffusion priors for 3D generation, feed-forward pipelines that prioritize speed and determinism, and emerging native 3D generative models that attempt to learn 3D distributions directly.
@@
Takeaway: Reformulating 3D models using the 3D encoder and projector aligns the 3D space with the semantic space, enabling llms to understand spatial features. Complex Data Understanding (O1): Interpret point clouds and meshes using 3D encoders and projection adapters hong20233d ; xiong20253ur , which fuses 3D spatial features with natural language semantics. NL-Based Interface (O2): LLMs function as high-capacity planners huang2024protchat , decomposing tasks and orchestrating specialized models to interact with natural language. Semantic Operators (O3): LLMs perform geometric relation identification and structured reasoning via methods such as Geometry CoT wang2025large , which translates complex 3D structures into textual descriptions wang20253dsmiles . Autonomous Evolution (O4):Learn 3D distributions directly bypassing 2D priors, and use techniques like mesh codebooks chen2024meshanything to achieve inherently 3D-consistent content creation.
@@ 5 Heterogeneous Data Analysis
XMODE nooralahzadeh2024explainable assumes that heterogeneous metadata is organized within a database and leverages LLMs to decompose natural language queries into subtasks such as text-to-SQL generation and image analysis. For instance, given the query “Plot the number of paintings that depict war in each century”, XMODE generates an operation sequence comprising: $(i)$ retrieving painting metadata from the database via SQL, $(ii)$ applying VLMs to determine whether each image depicts war, and $(iii)$ producing visualization code. In practice, however, heterogeneous data is often disorganized. To address this challenge, Wang et al. wang2024interactive encode data objects using modality-specific models (e.g., BERT for text, CLIP for images) to construct an index. They further introduce a vector weight learning model that adaptively adjusts modality weights to better capture their relative importance in similarity measurement wang2024must . The user query is used to retrieve relevant data through semantic similarity for final answer generation.
Takeaway: Heterogeneous data analysis agents exhibit strong capabilities in the following aspects: Complex Data Understanding (O1): Capture both structural and semantic information across diverse modalities. NL-Based Interface (O2): Facilitate cross-modal alignment, retrieval, and reasoning through natural language interaction. Semantic Operators (O3): Combine LLM-driven semantic operations with cross-modal processing capabilities. Autonomous Evolution (O4): Enable self-designing and adaptive evolution of analysis pipelines in response to changes in data sources.
@@ 6.1 Structured Data
Complex Analysis. Structured data encapsulates both semantic information (e.g., cell contents in tables) and structural information (e.g., key–value relationships in tables or edge connections in graphs). Effective analysis requires integrating these two dimensions. For simple tasks, LLMs can often derive answers directly from explicit semantic cues or straightforward structural patterns. However, complex tasks demand deeper comprehension and multi-hop reasoning across semantics and structure, which remain challenging for current LLMs. When LLMs alone fall short in handling such tasks end-to-end, multi-agent collaboration emerges as a promising direction. By decomposing a complex task into simpler subtasks, each agent can specialize in a narrower scope. Through prompt engineering or fine-tuning, LLMs can be tailored for specific subtasks, while intermediate outputs enhance both performance and interpretability.
Open World Task Adaptation. Structured data spans diverse domains with varying specifications and task requirements. LLMs trained for general-purpose analysis may fail to meet domain-specific needs. For instance, analyzing daily forms differs substantially from rigorous analyses in business or climate applications. Moreover, heterogeneity in system specifications, such as SQL dialects, poses additional adaptation challenges. To address these issues, domain adaptation through task-specific training or fine-tuning is essential. As manual annotation is often insufficient to support large-scale training, efficient automatic data generation methods warrant further exploration. In addition, techniques such as automatic query conversion and intermediate representations are crucial for bridging modality gaps, particularly between natural language and diverse SQL dialects.